The 8 acts for a talent show are being ordered randomly. What is the probability that the dance group is first followed by the comedian?


1
Expert's answer
2021-01-29T05:17:12-0500

The specified event consists of two events that must occur one after the other:

1. A dance group must be selected from eight participants. The probability of this event:


"P_1=1\/8."


2. A comedian must be selected from the seven remaining participants. The probability of this event


"P_2=1\/7."


The required probability value is equal to the product of the probabilities of these two events:


"P=P_1\cdot P_2=\frac{1}{8}\cdot \frac{1}{7}=0.018."


Answer: 0.018 or 1.8 %.